OpenStudy (anonymous):

The relation between time and distance is distance(s)=3t+2t^2. Find Acceleration.

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

you are given distance s=3t+t^2 rate of change of distance traveled is speed and rate of change of velocity is acceleration ds/dt = ds is the distance traveled in time dt ds/dt= rate of change of s wrt time t....later on u'll come to know abt differentiation here we are differentiating S wrt time t.. suppose y= x^n then dy/dx= n x^(n-1) u r being given... s=3t+t^2 v= ds/dt= 3 t^(1-1) + 2 t^(2-1) = 3+2t so v= ds/dt

OpenStudy (anonymous):

again rate of change of velocity is acceleration....i.e. dv/dt=a dv is the change in velocity in time dt so u have v=3+2t a= 0+ 2*t^(1-1)=2 so u got constant acceleration, which does not change with time.

OpenStudy (ghazi):

\[\frac{ dS }{ dt }= V(velocity)=\frac{ d(3t+2t^2) }{ dt }=3+4t\] now \[\frac{ dV }{ dt }= a(acceleration)=\frac{ d(3+4t) }{ dt }= 4\]

OpenStudy (anonymous):

sorry i wrote s=3 t +t^2 but u can solve for s= 3t +2 t^2

OpenStudy (amistre64):

\[\Delta x=v_ot+\frac 12at^2\]\[\frac 12 a=2\] \[a=4\]

OpenStudy (anonymous):

The relation between time and distance is distance(s)=3t+2t^2. Find Acceleration. v = d(3t+2t^2)/dx v = 3+4t a = d(3+4t)/dx a = 0 + 4 a = 4